### v2

@@ 4.2 Distillation Expands the Reasoning Boundary
Figure 7: Coverage comparison of base, Instruct, RLVR, and distilled models.
In addition to direct RL training, another effective approach to improving the reasoning ability of small base models is distillation from a powerful reasoning model Guo et al. (2025). This process is analogous to instruction-following fine-tuning in post-training. However, instead of using short instruction-response pairs, the training data consist of long CoT reasoning traces generated by the teacher model. Given the limitations of current RLVR in expanding reasoning capabilities, it is natural to ask whether distillation exhibits similar behavior. We focus on a representative model, DeepSeek-R1-Distill-Qwen-7B, which distills DeepSeek-R1 into Qwen2.5-Math-7B. We compare it with the base model Qwen2.5-Math-7B and its RL-trained counterpart Qwen2.5-Math-7B-Oat-Zero and include Qwen2.5-Math-7B-Instruct as an additional baseline. As shown in Figure 7, the pass@$k$ curve of the distilled model is consistently and significantly above that of the base model. This indicates that, unlike RL that is fundamentally bounded by the reasoning capacity of the base model, distillation introduces new reasoning patterns learned from a stronger teacher model. As a result, the distilled model is capable of surpassing the reasoning boundary of the base model.
Figure 8: Different RL algorithms. We use a folded y-axis range to better highlight the details at $k=1$ and $256$. Unfolded version can be found in Figure 14. The detailed values for each point at pass@1 and pass@256 are provided in Table 2 and Table 3.
@@ 4.3 Effects of Different RL Algorithms
Experiment Setup. We re-implement popular RL algorithms using the VeRL framework Sheng et al. (2024) for fair comparison, including PPO Schulman et al. (2017), GRPO Shao et al. (2024), Reinforce++ Hu (2025), RLOO Ahmadian et al. (2024), ReMax Li et al. (2024), and DAPO Yu et al. (2025). Following DAPO Yu et al. (2025) and Oat-Zero Liu et al. (2025b), we remove the KL term to avoid constraining model learning. During training, we use a prompt batch size of 128 and sample 8 responses per prompt. We set the maximum rollout length to 8,192 tokens and set the temperature to 1.0.
@@
Results. As shown in Figure 8 (top), although different RL algorithms exhibit slight variations in both pass@1 and pass@256, these differences are not fundamental. Different RL algorithms yield slightly different $\Delta_{\text{SE}}$ values (i.e., ranging from GRPO’s 43.9 to RLOO’s best 42.6 on the in-domain test set). Furthermore, we observe that $\Delta_{\text{SE}}$ remains consistently above 40 points across different algorithms, highlighting that existing RL methods are still far from achieving optimal sampling efficiency. This suggests that novel RL algorithms or entirely new paradigms may be necessary to approach the upper bound. Additional observations can be found at Section D.4.
@@ 4.4 Effects of RL Training
Asymptotic Effects. Based on the setup in Section 4.3, we investigate the effect of the training steps on the asymptotic performance of the model. As shown in Figure 1 (right), as RL training progresses, pass@1 on the training set consistently improves from 26.1 to 42.5. However, as RLVR training progresses, pass@256 progressively decreases, indicating a reduced reasoning boundary.
Effect of Number of Rollouts $n$. The training hyperparameter $n$, the number of responses per prompt, can affect pass@$k$ by enabling broader exploration during training. We increase $n$ from 8 to 32. As shown in Figure 15, pass@$k$ improves slightly over $n=8$, but the RL-trained model is still eventually outperformed by the base model. We leave the question of whether scaling RLVR training can eventually surpass the base model to future investigation.
Effect of KL Loss. To control model deviation, some prior work adds a KL penalty. We ablate this by applying a KL term with coefficient 0.001. As shown in Figure 15, the KL-regularized model achieves similar pass@1 to GRPO without KL, but with a much lower pass@128.
@@ 4.5 Effects of Entropy
As RL training progresses, the model’s output entropy typically decreases Yu et al. (2025), which may contribute to a reduced reasoning boundary due to less diverse output. To investigate this factor, we increase the generation temperature of the RLVR-trained model to match the output entropy of the base model at $T=0.6$. As shown in Figure 17, although the RLVR model performs slightly better pass@$k$ at higher temperatures compared to its own performance at $T=0.6$, it still underperforms the base model across pass@$k$. This suggests that while reduced entropy contributes to the narrowing of the reasoning boundary, it alone does not fully account for the reduction.
@@ 6 Conclusion and Discussion
RLVR is widely regarded as a promising approach to enable LLMs to continuously self-improve and acquire novel reasoning capabilities. In this paper, we systematically investigate the effect of current RLVR methods on the reasoning capacity boundaries of LLMs. Surprisingly, our findings reveal that current RLVR does not elicit fundamentally new reasoning patterns; instead, the reasoning capabilities of RLVR-trained models remain bounded by those of their base models. These results indicate that current RLVR methods have not fully realized the potential of reinforcement learning to elicit novel reasoning abilities in LLMs through exploration and exploitation. This limitation may stem from the lack of effective exploration strategies in the vast language space or the absence of multi-turn agent-environment interactions needed to generate novel experience. We provide a more detailed discussion of the possible causes of this gap and promising future directions in Appendix C.
@@ A.2 Low-Variance pass@$k$ Estimation
[…]
@@
In our experiments, we set $n$ to the largest (i.e., rightmost) $k$ value in the pass@$k$ curves, typically 128, 256, or 1024. For example, in Figure 2, we use $n=128$ for MATH500, Minerva, and GSM8K, and $n=1024$ for AMC23 and AIME24. For the Olympiad benchmark, we set $n=128$ for the Qwen models and $n=1024$ for LLaMA-3.1-8B, due to its relatively lower base model capacity.
@@ Appendix C Discussion
In Section 3 and Section 4, we identified key limitations of RLVR in enhancing LLM reasoning capabilities. In this section, we explore possible underlying factors that may explain why RLVR remains bounded by the reasoning capacity of the base model.
@@
Discussion 2: Priors as a Double-Edged Sword in This Vast Action Space. Since the sampling of responses is guided by the pretrained prior, the policy may struggle to explore new reasoning patterns beyond what the prior already provides. Specifically, in such a complex and highly combinatorial space, most responses generated during training are constrained by the base model’s prior. Any sample deviating from the prior is highly likely to produce invalid or non-sensical outputs, leading to negative reward. As discussed in Section 2.1, policy gradient algorithms aim to maximize the log-likelihood of responses within the prior that receive positive rewards, while minimizing the likelihood of responses outside the prior that receive negative rewards. As a result, the trained policy tends to produce responses already present in the prior, constraining its reasoning ability within the boundaries of the base model. From this perspective, training RL models from a distilled model may temporarily provide a beneficial solution, as distillation helps inject a better prior.
@@ D.2 Validity of Chain-of-Thought on AIME24
The base and RL model pass@$k$ curves on this filtered AIME24 can be found in Figure 12, showing a similar trending to previous results. Although this filtering method is heuristic, it proves to be effective. Applying it to AIME24 (30 questions) results in a subset of 18 questions. We then prompt the models to answer these filtered questions using CoT reasoning. Then we perform a manual inspection of all CoTs that led to correct answers on the hardest problems – those with an average accuracy below 5%. The base model answered 7 such questions, with 5 out of 6 containing at least one correct CoT (excluding one ambiguous case of correctness due to skipped reasoning steps). Similarly, the RL-trained model answered 6 questions, 4 of which included at least one correct CoT. These results suggest that even for the hardest questions in the challenging AIME24, base model can sample valid reasoning paths to solve the problems.
@@ D.3 Accuracy Distribution Visulization
D.4 Different RLVR Algorithms
We report several additional observations on different RLVR algorithms in Figure 8. First, DAPO achieves slightly higher pass@1 scores across all three datasets; however, its dynamic sampling strategy requires approximately 36×3$\sim$ 6$\times$3 ∼ 6 × more samples per batch during training compared to other algorithms. Moreover, its performance drops significantly at $k=256$. Second, RLOO and Reinforce++ perform consistently well across the entire $k$ range (from 1 to 256), while maintaining efficient training costs, achieving a good balance between effectiveness and efficiency. Third, ReMax shows lower performance at both pass@1 and pass@256. We hypothesize that this is due to its use of the greedy response reward as the advantage baseline, which in the RLVR setting is binary (0 or 1) and highly variable. This likely results in unstable gradient updates during training.
Figure 14: Unfolded y-axis version of Figure 8.
Table 2: Detailed values for each point at pass@1 and pass@256 across different RL algorithms in Figure 8.
@@ D.4 Different RLVR Algorithms
Table 3: Detailed values at pass@1 and pass@256 across different RL training steps in Figure 1 (right).
@@
D.5 Effects of KL and Rollout Number
@@
D.6 Solvable Problem Coverage Analysis
@@
Qwen2.5-7B-Instruct-1M 400, 402, 403, 407, 409, 412, 413, 417, 418, 419, 422, 423, 427, 432, 433, 436, 438, 439, 440, 444, 445, 448, 449
Coder-R1 400, 402, 403, 407, 412, 413, 417, 418, 419, 422, 423, 427, 430, 433, 438, 439, 440, 444, 445, 449
D.7 Temperature and Entropy Analysis
@@
D.8 Training Dynamics
Figure 18: The curves of training reward, response length, and generation entropy during training, corresponding to experiments in Section 4.
D.9 CoT Case Analysis
@@ Appendix E Prompt Templates
We provide the prompt templates used for training and evaluation in our experiments. The prompt for SimpleRL training and evaluation is shown in Figure 21, while the prompt for Oat-Zero is shown in Figure 22. For Code-R1 training, prompt in Figure 23 is adopted. For Code-R1 evaluation, we follow the original codebase and adopt the default templates from the benchmarks, including LiveCodeBench prompt (Figure 24), HumanEval+, and MBPP+ prompt (Figure 25). The prompt used for EasyR1 training and evaluation is shown in Figure 26. For VeRL-trained RL models, as discussed in Section 4.3 and Section 4.4, the training and evaluation prompts are provided in Figure 27. To ensure a fair comparison, the base models use the same prompts as their corresponding RL-trained counterparts during evaluation.
@@
Figure 24: Since Code-R1 does not specify an evaluation prompt, we adopt the original LiveCodeBench evaluation prompt. To encourage both the base and RL-trained models to generate code, we append ‘‘‘python to the end of the prompt. Using this setup, we reproduce a pass@$1$ score of 28.6, which is close to the reported 29.7.
@@
Figure 27: (1) Prompt for VeRL training on Omni-math-train and evaluation on Omni-math-eval and MATH500.

### v3

@@ 4.2 Distillation Expands the Reasoning Boundary
Figure 7: pass@$k$ of base, Instruct, RLVR, and distilled models.
In addition to direct RL training, another effective approach to improving the reasoning ability of small base models is distillation from a powerful reasoning model Guo et al. (2025). This process is analogous to instruction-following fine-tuning in post-training. However, instead of using short instruction-response pairs, the training data consist of long CoT reasoning traces generated by the teacher model. Given the limitations of current RLVR in expanding reasoning capabilities, it is natural to ask whether distillation exhibits similar behavior. We focus on a representative model, DeepSeek-R1-Distill-Qwen-7B, which distills DeepSeek-R1 into Qwen2.5-Math-7B. We compare it with the base model Qwen2.5-Math-7B and its RL-trained counterpart Qwen2.5-Math-7B-Oat-Zero and include Qwen2.5-Math-7B-Instruct as an additional baseline. As shown in Figure˜7, the pass@$k$ curve of the distilled model is consistently and significantly above that of the base model. This indicates that, unlike RL that is fundamentally bounded by the reasoning capacity of the base model, distillation introduces new reasoning patterns learned from a stronger teacher model. As a result, the distilled model is capable of surpassing the reasoning boundary of the base model.
@@ 4.3 Effects of Different RL Algorithms
Experiment Setup. We re-implement popular RL algorithms using the VeRL framework Sheng et al. (2024) for fair comparison, including PPO Schulman et al. (2017), GRPO Shao et al. (2024), Reinforce++ Hu (2025), RLOO Ahmadian et al. (2024), ReMax Li et al. (2024), and DAPO Yu et al. (2025). Following DAPO Yu et al. (2025) and Oat-Zero Liu et al. (2025b), we remove the KL term to avoid constraining model learning. During training, we use the AdamW optimizer Loshchilov & Hutter (2017) with a constant learning rate of $10^{-6}$. For rollout, we employ a prompt batch size of 256 and generate 8 responses per prompt. The maximum rollout length is set to 8,192 tokens, and the sampling temperature is set as 1.0. We use a PPO mini-batch size of 256.
@@
Results. As shown in Figure˜8 (top), although different RL algorithms exhibit slight variations in both pass@1 and pass@256, these differences are not fundamental. Different RL algorithms yield slightly different $\Delta_{\text{SE}}$ values (i.e., ranging from GRPO’s 43.9 to RLOO’s best 42.6 on the in-domain test set). Furthermore, we observe that $\Delta_{\text{SE}}$ remains consistently above 40 points across different algorithms, highlighting that existing RL methods are still far from achieving optimal sampling efficiency. This suggests that novel RL algorithms or entirely new paradigms may be necessary to approach the upper bound. Additional observations can be found at Section˜D.5.
Figure 8: (Top) Different RL algorithms. (Bottom) Different RL training steps. The detailed values for each point at pass@1 and pass@256 are provided in Table˜3 and Table˜4.
@@ 4.4 Effects of RL Training
Asymptotic Effects. Based on the setup in Section˜4.3, we investigate the effect of the training steps on the asymptotic performance of the model. As shown in Figure˜1 (right), as RL training progresses, pass@1 on the training set consistently improves from 26.1 to 42.5. However, as RLVR training progresses, pass@256 progressively decreases, indicating a reduced reasoning boundary.
Effect of Number of Rollouts $n$. The training hyperparameter $n$, the number of responses per prompt, can affect pass@$k$ by enabling broader exploration during training. We increase $n$ from 8 to 32. As shown in Figure˜16, pass@$k$ improves slightly over $n=8$, but the RL-trained model is still eventually outperformed by the base model. We leave the question of whether scaling RLVR training can eventually surpass the base model to future investigation.
Effect of KL Loss. To control model deviation, some prior work adds a KL penalty. We ablate this by applying a KL term with coefficient 0.001. As shown in Figure˜16, the KL-regularized model achieves similar pass@1 to GRPO without KL, but with a much lower pass@128.
@@ 4.5 Effects of Entropy
As RL training progresses, the model’s output entropy typically decreases Yu et al. (2025), which may contribute to a reduced reasoning boundary due to less diverse output. To investigate this factor, we increase the generation temperature of the RLVR-trained model to match the output entropy of the base model at $T=0.6$. As shown in Figure˜18, although the RLVR model performs slightly better pass@$k$ at higher temperatures compared to its own performance at $T=0.6$, it still underperforms the base model across pass@$k$. This suggests that while reduced entropy contributes to the narrowing of the reasoning boundary, it alone does not fully account for the reduction.
4.6 Effects of Model Size Scaling
Scaling plays a central role in the capabilities of contemporary LLMs. It remains an important question whether the conclusions drawn continue to hold as model size increases. For many large models, isolating the effect of RLVR is not feasible. For example, in the case of GPT-o1, the base model is not publicly accessible. Qwen3-235B Yang et al. (2025a) is trained through multiple stages, including RLVR and long-context CoT supervised fine-tuning, which makes it impossible to disentangle the impact of RLVR alone. For Deepseek-R1-Zero, the absence of a publicly hosted API forced us to self-host the model, but throughput was limited to around 50 tokens per second at a maximum sequence length of 32k, rendering pass@$k$ evaluation currently impractical. As a more tractable alternative, we selected the Magistral-Medium-2506 API to conduct a preliminary set of experiments. This model is trained using pure RL, with Mistral-Medium-3-2505 as the starting model Rastogi et al. (2025). Although the model size is not disclosed, Magistral-Medium performs comparably to Deepseek-R1 and is positioned near the frontier in terms of reasoning capability.
Figure 9: pass@$k$ curves of Magistral-Medium.
We queried the models using a maximum context length of 40k as the original paper does. Once again, we observed that RLVR provides significant gains at low $k$, but little or no improvement at higher $k$. Specifically, at $k=1$, the RLVR-enhanced model solves approximately 7 more problems on AIME24 and 8 more on AIME25 compared to its base version. However, as $k$ increases, the performance gap steadily narrows. These observations suggest that our conclusion continues to hold even for current, highly capable, near-frontier reasoning models. Whether this trend persists as more compute, such as pre-training scale budgets, is dedicated to RL training remains a critical question for the future of LLM reasoning.
@@ 6 Conclusion and Discussion
RLVR is widely regarded as a promising approach to enable LLMs to continuously self-improve and acquire novel reasoning capabilities. In this paper, we systematically investigate the effect of current RLVR methods on the reasoning capacity boundaries of LLMs. Surprisingly, our findings reveal that current RLVR does not elicit fundamentally new reasoning patterns; instead, the reasoning capabilities of RLVR-trained models remain bounded by those of their base models. These results indicate that current RLVR methods have not fully realized the potential of reinforcement learning to elicit novel reasoning abilities in LLMs through exploration and exploitation. This limitation may stem from the lack of effective exploration strategies in the vast language space or the absence of multi-turn agent-environment interactions needed to generate novel experience. We provide a more detailed discussion of the possible causes of this gap and promising future directions in Appendix˜C.
7 Acknowledgements
This work is supported in part by the National Key R&D Program of China under Grant 2022ZD0114903, the National Natural Science Foundation of China under Grants 42327901 and U24B20173, and the Scientific Research Innovation Capability Support Project for Young Faculty under Grant ZYGXQNJSKYCXNLZCXM-I20.
@@ A.2 Low-Variance pass@$k$ Estimation
$\text{pass@}k:=\mathbb{E}_{x_{i}\sim\mathcal{D}}\left[1-\frac{\binom{n-c_{i}}{k}}{\binom{n}{k}}\right$] (2)
@@
In our experiments, we set $n$ to the largest (i.e., rightmost) $k$ value in the pass@$k$ curves, typically 128, 256, or 1024. For example, in Figure˜2, we use $n=128$ for MATH500, Minerva, and GSM8K, and $n=1024$ for AMC23 and AIME24. For the Olympiad benchmark, we set $n=128$ for the Qwen models and $n=1024$ for LLaMA-3.1-8B, due to its relatively lower base model capacity.
@@ Appendix C Discussion
In Section˜3 and Section˜4, we identified key limitations of RLVR in enhancing LLM reasoning capabilities. In this section, we explore possible underlying factors that may explain why RLVR remains bounded by the reasoning capacity of the base model.
@@
Discussion 2: Priors as a Double-Edged Sword in This Vast Action Space. Since the sampling of responses is guided by the pretrained prior, the policy may struggle to explore new reasoning patterns beyond what the prior already provides. Specifically, in such a complex and highly combinatorial space, most responses generated during training are constrained by the base model’s prior. Any sample deviating from the prior is highly likely to produce invalid or non-sensical outputs, leading to negative reward. As discussed in Section˜2.1, policy gradient algorithms aim to maximize the log-likelihood of responses within the prior that receive positive rewards, while minimizing the likelihood of responses outside the prior that receive negative rewards. As a result, the trained policy tends to produce responses already present in the prior, constraining its reasoning ability within the boundaries of the base model. From this perspective, training RL models from a distilled model may temporarily provide a beneficial solution, as distillation helps inject a better prior.
@@ D.2 Validity of Chain-of-Thought on AIME24
The base and RL model pass@$k$ curves on this filtered AIME24 can be found in Figure˜13, showing a similar trending to previous results. Although this filtering method is heuristic, it proves to be effective. Applying it to AIME24 (30 questions) results in a subset of 18 questions. We then prompt the models to answer these filtered questions using CoT reasoning. Then we perform a manual inspection of all CoTs that led to correct answers on the hardest problems – those with an average accuracy below 5%. The base model answered 7 such questions, with 5 out of 6 containing at least one correct CoT (excluding one ambiguous case of correctness due to skipped reasoning steps). Similarly, the RL-trained model answered 6 questions, 4 of which included at least one correct CoT. These results suggest that even for the hardest questions in the challenging AIME24, base model can sample valid reasoning paths to solve the problems.
@@ D.3 Accuracy Distribution Visulization
D.4 Perplexity Analysis
To analyze how perplexity evolves over the course of RLVR training, we evaluated three RLVR checkpoints–early, middle, and final in Section˜4.3 RL training. For each checkpoint, we sampled 32 responses per problem, computed the median among 32 perplexity values, and reported the average over the first 10 problems in the table. As expected, we observed that $\text{PPL}_{\text{Base}}(\mathbf{Y}_{\text{RL}}|x)$ gradually decreases as RL training progresses, indicating that RLVR mainly sharpens the distribution within the base model’s prior rather than expanding beyond it.
Figure 15: Perplexity Evolution during RL Training.
D.5 Different RLVR Algorithms
We report several additional observations on different RLVR algorithms in Figure˜8. First, DAPO achieves slightly higher pass@1 scores across all three datasets; however, its dynamic sampling strategy requires approximately $3\sim 6\times$ more samples per batch during training compared to other algorithms. Moreover, its performance drops significantly at $k=256$. Second, RLOO and Reinforce++ perform consistently well across the entire $k$ range (from 1 to 256), while maintaining efficient training costs, achieving a good balance between effectiveness and efficiency. Third, ReMax shows lower performance at both pass@1 and pass@256. We hypothesize that this is due to its use of the greedy response reward as the advantage baseline, which in the RLVR setting is binary (0 or 1) and highly variable. This likely results in unstable gradient updates during training.
Table 3: Detailed values for each point at pass@1 and pass@256 across different RL algorithms in Figure˜8.
@@ D.5 Different RLVR Algorithms
Table 4: Detailed values at pass@1 and pass@256 across different RL training steps in Figure˜1 (right).
@@
D.6 Effects of KL and Rollout Number
@@
D.7 Solvable Problem Coverage Analysis
Table˜2 reports the fraction of problems categorized as four conditions: (1) both models solve the problem at least once, (2) only the base model solves it, (3) only the RLVR model solves it, and (4) neither model solves it in any of the $k$ samples. It highlights that there are many cases where the base model solves a problem but RLVR fails (type 2), and very few where RLVR succeeds while the base does not (type 3). Even in the rare type 3 cases (e.g., 1% or about 5 problems in MATH500), the base model is able to solve all of them when sampling 1024 times. These results support our conclusion that RLVR rarely solves problems the base model cannot and generally results in reduced coverage.
@@ D.7 Solvable Problem Coverage Analysis
Qwen2.5-7B-Instruct-1M
400, 402, 403, 407, 409, 412, 413, 417, 418, 419, 422, 423,
427, 432, 433, 436, 438, 439, 440, 444, 445, 448, 449
Coder-R1
400, 402, 403, 407, 412, 413, 417, 418, 419, 422, 423,
427, 430, 433, 438, 439, 440, 444, 445, 449
D.8 Temperature and Entropy Analysis
@@
D.9 Training Dynamics
Figure 19: The curves of training reward, response length, and generation entropy during training, corresponding to experiments in Section˜4.
D.10 CoT Case Analysis
@@ Appendix E Prompt Templates
We provide the prompt templates used for training and evaluation in our experiments. The prompt for SimpleRL training and evaluation is shown in Figure˜22, while the prompt for Oat-Zero is shown in Figure˜23. For Code-R1 training, prompt in Figure˜24 is adopted. For Code-R1 evaluation, we follow the original codebase and adopt the default templates from the benchmarks, including LiveCodeBench prompt (Figure˜25), HumanEval+, and MBPP+ prompt (Figure˜26). The prompt used for EasyR1 training and evaluation is shown in Figure˜27. For VeRL-trained RL models, as discussed in Section˜4.3 and Section˜4.4, the training and evaluation prompts are provided in Figure˜28. For evaluating Mistral and Magistral models on AIME24/25, prompts are provided in Figure˜29. To ensure a fair comparison, the base models use the same prompts as their corresponding RL-trained counterparts during evaluation.
@@
Figure 25: Since Code-R1 does not specify an evaluation prompt, we adopt the original LiveCodeBench evaluation prompt. To encourage both the base and RL-trained models to generate code, we append ‘‘‘python to the end of the prompt. Using this setup, we reproduce a pass@11 score of 28.6, which is close to the reported 29.7.
@@
Figure 28: Prompt for VeRL training on Omni-math-train and evaluation on Omni-math-eval and MATH500.
Figure 29: Prompt for Mistral and Magistral model evaluation.